2. A sports center sells a pair of running shoes at a markup of 20% of the selling price. Their markup on a specific brand is P350.


A. How much is the cost of the running shoes?


B. What is the selling price?


C. What is the rate of markup based on cost?



3. An appliance store cells electric kettle at a markup of 18% of the selling price. The store's profit of a particular model is P310.50.


A. For how much does a store sells kettle?


B. What was the cost of a kettle when it was bought by the store?


C. What is the rate of markup based on the cost?



4. An appliance center sells washing machine at a markup of 15% of the selling price. The store's markup on a certain model is P300.


A. What is the selling price of the washing machine?


B. What is the original cost of the washing machine?


C. What is the rate markup based on the cost?



5. A sungka is priced P999.75 at a certain toy store. If they used a markup of 35% of the cost, what is the markup as a percent of the selling price

1
Expert's answer
2021-11-08T15:58:49-0500

(2)

a)Cost of the running shoes


Cost=Selling price-Markup


=P1750-P350


=P1400


b)selling price:


="\frac{P350}{0.2}=P1750"


c) rate of markup based on cost


=(Mark up"\div" cost)"\times 100\%"


=(350"\div" 1400)"\times 100\%"


=25"\%"


3)

A. For how much does a store sells kettle?


Selling price ="\frac{P310.5}{0.18}=P1725"


B. What was the cost of a kettle when it was bought by the store?


Cost=Selling price-Markup


=P1725-P310.5


=P1414.5


C. What is the rate of markup based on the cost?


=(Mark up"\div" cost)"\times 100\%"


=(310.5"\div" 1414.5)"\times 100\%"


=21.95"\%"


4)

A. What is the selling price of the washing machine?


Selling price="\frac{P300}{0.15 }"


=P2000


B. What is the original cost of the washing machine?


Cost=Selling price-Markup


=2000-300


=P1700



C. What is the rate markup based on the cost?


=(Mark up"\div" cost)"\times 100\%"


=(300"\div" 1700)"\times 100\%"


=17.65"\%"




5. The markup as a percent of the selling price.


=((35"\%\div" ("35" "\%" "+1"))


="\frac{0.35}{0.35+1}\times 100"


="25.9\%"